OpenStudy (anonymous):

Find the range of values of k such that the circle \(x ^{2}+y ^{2}+2x-4y-13=0\) and the straight line x-y+k=0 intersect at two distinct points. A. -93 D. k<-3 or k>9 @Callisto @appleduardo

OpenStudy (callisto):

x = y-k <- Sub this into the equation of circle, you'll have an quadratic equation. 2 distinct intersections implies delta > 0, then you can solve for the range of k.

OpenStudy (callisto):

Maybe it's better to sub y=x+k...

OpenStudy (anonymous):

my final step is like this\[2y ^{2}-2y-2ky+k ^{2}-2k-13=0\] what's wrong with this step?? :O

OpenStudy (anonymous):

you know the slope of the given line this line can be at a max. of radius of given circle

OpenStudy (anonymous):

max. distance is from center

OpenStudy (callisto):

Can be, but how can you be sure?

OpenStudy (anonymous):

because lines makes two intercept with given circle

OpenStudy (anonymous):

@Callisto i got it -3<k<9 ?

OpenStudy (anonymous):